OpenStudy (anonymous):

Jesse kicks a soccer ball off the ground and in the air with an initial velocity of 29 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (anonymous):

H(t) = -16t^2 + vt + s H'(t) = -32t + v H'(0) = -32(0) + v = 29 v = 29 At maximum height, H'(t) = 0. -32t + 29 = 0 -32t = -29 t = 29 / 32 H(29 / 32) = -16(29 / 32)^2 + 29(29 / 32) + s = 841 / 64 + s

OpenStudy (anonymous):

I don't understand how to solve that equation.

OpenStudy (anonymous):

s should be the initial height, which is 0 since the ball is kicked off the ground

OpenStudy (anonymous):

So, just divide 841 by 64?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Alright, thank you.